A whispering gallery has a semislliptical chilling that is 10m high and 32m. How high is the ceiling above the two foci?
1
Expert's answer
2020-11-17T10:37:33-0500

"\text{We have the equation of an ellipse with semiaxes a= 32 and b =10}"

"\frac{x^2}{a^2}+\frac{y^2}{b^2}=1; a^2=b^2+c^2"

"\text {where }c \text{ half the focal length. }"

"\text{for }x =c \text{ value }y \text{ is the desired value}"

"y^2= b^2(1-\frac{x^2}{a^2})"

"\text {for } x=c;\ y^2=b^2(1-\frac{a^2-b^2}{a^2})=\frac{b^4}{a^2}"

"y=\frac{b^2}{a}=3.125"

Answer: 3.125 m